OpenStudy (anonymous):

A box of 15 parts contains 7 that are defective. A worker picks parts one at a time and attempts to install them. Find the probability of each outcome in​ Find the probability of each outcome in b) At least one of the first three is good. c) The first four picked are all good. d) The worker has to pick five parts to find that is good. (Part d is asking for the probability that the first four parts are bad, and the fifth part is good.

OpenStudy (anonymous):

b) At least one of the first three is good. c) The first four picked are all good. d) The worker has to pick five parts to find that is good. (Part d is asking for the probability that the first four parts are bad, and the fifth part is good.

OpenStudy (anonymous):

pl help me out im so lost on it

OpenStudy (anonymous):

A box of 15 parts contains 7 that are defective. A worker picks parts one at a time and attempts to install them. Find the probability of each outcome in​ Find the probability of each outcome in

OpenStudy (anonymous):

STop SPAMMING ME

OpenStudy (kropot72):

This question involves sampling without replacement. To solve "b) At least one of the first three is good. " we can first find the probability that there is zero good in the first three. Then we subtract that probability from 1 to find the required probability: \[\large P(zero\ good)=\frac{7C3}{15C3}=\frac{1}{13}\] \[\large P(at\ least\ 1\ good)=\frac{1}{1}-\frac{1}{13}=you\ can\ calculate\]

OpenStudy (kropot72):

@vampyradeatwish Are you there?